# Fluctuation Theorem on Riemannian Manifold

### 29 Dec 2023 v1

@@ I Introduction
In the past twenty five years or so, there have been some repeated attempts in the construction of LE on Riemannian manifolds in the mathematical literature [14, 15, 16, 17, 18, 19]. However, the first law and other thermodynamic relations were not considered in these works. Xing et al. [20, 21] explored an Ito-type nonlinear LE and established a FT on Riemannian manifold. It will be clear in the following main text that the meaning of “covariance” in their works is totally different from ours.
@@ II.1 Underdamped case
In [22], the relativistic covariant underdamped Langevin equation (ULE) on pseudo-Riemannian spacetime is established. The same procedure can be used for constructing LE on Riemannian space $M$ with a positive definite metric $g_{\mu\nu}(x)$, so we directly present the result,
@@
where $R^{\mu}{}_{a}$ represent the stochastic amplitudes which may depend on $\tilde{x}^{\mu}$ and $\tilde{p}^{\mu}$, and ${\rm d}\tilde{w}^{a}_{t}$ are Gaussian noises with probability distribution functions (PDFs)
@@
The symbol $\circ_{S}$ represents the Stratonovich coupling which ensures the chain rule in stochastic calculus. Greek indices label spatial directions and Latin indices label independent noises, which are all running from 1 to $d$. Tilded variables like ${\rm d}\tilde{w}_{t}^{a}$ represent random variables, and the un-tilded symbols like ${\rm d}w^{a}$ their realizations.
@@
(1) The LE described above is a system of stochastic differential equations on the tangent bundle $TM$ regarded as the space of micro states of the BP. Recall that $TM$ is equipped with the Sasaki metric [23]
@@
(2) Our approach to LE is the traditional one, as opposed to the more abstract nonlinear approach used in [20, 21]. Eq. (2) can be viewed as the geodesic equation supplemented by additional force terms, including a stochastic force
@@
a damping force $f^{\mu}_{\text{dp}}:=-K^{\mu}{}_{\nu}\tilde{p}_{t}^{\nu}/m$ and an external force $f^{\mu}_{\text{ex}}$. The second term in the stochastic force is known as additional stochastic force [24, 25], which is required in order for the BP to be able to reach thermal equilibrium with the heat reservoir. In one-dimensional case, the stochastic force can also be expressed in the form of post-point rule $\xi=R\circ_{p}{\rm d}\tilde{w}_{t}/{\rm d}t$, hence some authors argued that the post-point rule is better suited for LE with multiplicative noises. However, in higher-dimensional cases, the post-point rule leads to a different result,
@@
(3) The use of Stratonovich coupling ensures the covariance in the usual sense in Riemannian geometry. This is different from previous works [16, 17, 18, 19] which adopted Ito coupling and interpreted the covariance in the sense of jet bundle geometry.
(4) Although eqs. (1)-(2) look the same as their relativistic counterparts [22], there are some essential differences. First, the time $t$ used here is absolute, meaning that eqs. (1)-(2) are non-relativistic; Second, the heat reservoir hiding behind the stochastic and damping force terms is also non-relativistic, there is no need to worry about the relativistic effects such as the Tolman-Ehrenfest red shift; Third, the momentum space is flat, as opposed to the relativistic case; Last, the damping coefficient $K^{\mu}{}_{\nu}$ is now assumed to be invertible, which is only possible when the number of independent noises equals the dimension $d$ of the manifold $M$.
The external force term depends on the position $x$ of the BP and an external control parameter $\lambda$, and can be separated into conservative and non-conservative parts,
@@
of the BP. In the extremal case with $\hat{f}^{\mu}=-f^{\mu}_{\text{con}}$, The energy will be consisted purely of the kinematic energy.
@@
where $\tilde{E}_{t}$ is the energy considered as a random variable and $E$ its realization. The part of the increase of energy caused by the heat reservoir is comprehended as trajectory heat ${\rm d}\tilde{Q}_{t}$, and the rest part as trajectory work ${\rm d}\tilde{W}_{t}$,
@@
It should be reminded that the ambiguity in the decomposition of the external force also leads to an ambiguity in the trajectory work $\tilde{W}_{t}$. However, in any case, the trajectory heat ${\rm d}\tilde{Q}_{t}={\rm d}\tilde{E}_{t}-{\rm d}\tilde{W}_{t}$ is always unambiguously defined.
The Fokker-Planck equation (FPE) associated to eqs. (1)-(2) can also be established on $TM$. The PDF for the BP under the measure ${\rm d}^{2d}X$ is denoted as $\Phi_{t}(x,p):=\Pr[\tilde{x}_{t}=x,\tilde{p}_{t}=p]$, and is clearly coordinate-independent. By use of the diffusion operator method [26], one can get
@@
is the Liouville vector field. Notice that, as long as only the PDF is concerned, there is also some ambiguity in the choice of $R^{\mu}{}_{a}$, because only the diffusion tensor $D^{\mu\nu}$ appears in the FPE (6).
Let the non-conservative force be temporarily turned off and the external control parameter be fixed. Then, after sufficiently long period of time, the BP will reach a thermal equilibrium with the heat reservoir, yielding the equilibrium PDF
$\Phi_{t}(x,p)=\frac{1}{Z}\exp\left[-\frac{1}{T}\left(\frac{g_{\mu\nu}p^{\mu}p^{\nu}}{2m}+U(x,\lambda)\right)\right].$
@@
$D^{\mu\nu}=2TK^{\mu\nu}.$ (7)
To facilitate the discussion about FT, we introduce the time-reversal transform (TRT)
$I:(x,p)\mapsto(x,-p).$
It is obvious that TRT preserves the metric, i.e. $I^{*}\hat{g}=\hat{g}$. Since the damping force $f^{\mu}_{\text{dp}}|_{X}=-K^{\mu}{}_{\nu}p^{\nu}/m$ reverses sign under TRT, the damping tensor must be invariant under TRT,
@@
Eq. (7) implies that the diffusion tensor is also invariant under TRT. As mentioned above, there is some freedom in choosing the stochastic amplitudes, and hence also in determining their behaviors under TRT. Here we assume the simplest transformation rule,
@@ II.2 Overdamped case
The stochastic mechanics characterized by LE is a branch of physics with multiple time scales, of which the smallest one is the time scale ${\rm d}t$ which allows for a sufficient number of collisions between the BP and the heat reservoir particles which cause little changes in the state of the BP [27]. If the temporal resolution $\Delta t$ greatly exceeds ${\rm d}t$ but is still smaller than the relaxation time, the ULE (1)-(2) emerges. If $\Delta t$ is further increased so that it is much larger than the relaxation time, the overdamped limit emerges.
The relaxation process of the BP can be viewed either as the process by which the damping force attains a state of mechanical equilibrium with other forces (mechanical relaxation process), or as the process by which the BP achieves local thermal equilibrium with the heat reservoir (thermodynamic relaxation process). The characteristic timescales associated with both processes are $m/\kappa$, where $\kappa$ is the eigenvalue of $K^{\mu\nu}$. Consequently, the OLE should arise when $\Delta t$ greatly exceeds $m/\kappa$.
@@
This result is achieved in flat space under the condition that the damping coefficients are position- and momentum-independent [28, 29], and is often viewed as a stochastic differential equation in configuration space. However, if the stochastic amplitudes are momentum-dependent, so is the additional stochastic force. Thus the mechanical equilibrium condition cannot be understood as a stochastic differential equation in configuration space. To avoid the above difficulty, let us consider the simpler situation in which the stochastic amplitudes are momentum-independent. Using the thermodynamic relaxation approach [30], it will be shown that, even in this simpler situation, a nontrivial additional stochastic force term still arises in the corresponding OLE.
@@
$\rho_{t}(x)=\frac{1}{Z}{\rm e}^{-U(x,\lambda)/T}$
is a solution of eq. (11) provided the non-conservative force is turned off and the external parameter is fixed.
@@
$\hat{R}^{\mu}{}_{a}:=\hat{K}^{\mu}{}_{\nu}R^{\nu}{}_{a},\,\hat{F}^{\mu}=\hat{K}^{\mu}{}_{\nu}f^{\nu}_{\text{ex}}+\frac{1}{2}\hat{R}^{\mu}{}_{a}\nabla_{\nu}\hat{R}^{\nu}{}_{a},$
@@
Since the inertial effect can be ignored in the overdamped case, the energy of the BP contains only the potential energy, i.e. $E(x,\lambda)=U(x,\lambda)$. Using eq. (9) and the chain rule, we have
@@ III Fluctuation theorem
Now we come to the stage for describing FT on Riemannian manifold based on the description of stochastic trajectories. Since the trajectory probability in continuous time is hard to deal with, we first take a discrete equidistant set of time points $t_{I}=t_{0}<t_{1}<...<t_{n}=t_{F}$ and then take the continuum limit at the end of the calculation.
@@
Before proceeding, it is necessary to clarify the concepts of ensemble and trajectory entropy productions. At the time $t$, the ensemble entropy of the BP reads
@@
and the ensemble entropy production in the process is $\Delta S=S_{t_{F}}-S_{t_{I}}$. Notice that
@@
is not the entropy of the overdamped BP, however, the ensemble entropy production of the BP can be represented as the difference of $-\overline{\ln\rho_{t}}$, because the ensemble entropy for the overdamped BP can be evaluated to be
@@
subtracting the initial value from the final value leaves only the difference of $-\overline{\ln\rho_{t}}$,
@@
Since the Brownian motion is a Markov process, the trajectory probability can be decomposed into the product of transition probabilities,
$\Pr[\tilde{X}_{[t]}=X_{[t]}]$ $=\prod_{i=0}^{n-1}\Pr[\tilde{X}_{i+1}=X_{i+1}|\tilde{X}_{i}=X_{i}]\Pr[\tilde{X}_{0}=X_{0}].$
@@
where, due to the Clausius equality, the change of the entropy of the heat reservoir is
@@
${\rm e}^{-\left\langle\Sigma_{\tilde{X}_{[t]}}\right\rangle}$ $\leq\left\langle{\rm e}^{-\Sigma_{\tilde{X}_{[t]}}}\right\rangle=\int\mathcal{D}[X_{[t]}]\Pr[\tilde{X}^{-}_{[t]}=X^{-}_{[t]}]$
@@
where $\mathcal{D}[X_{[t]}]={\rm d}X_{0}\wedge{\rm d}X_{1}\wedge...\wedge{\rm d}X_{n}$ is the measure in the trajectory space. It is easy to prove that the map $X_{[t]}\mapsto X^{-}_{[t]}$ preserves the measure, i.e. $\mathcal{D}[X_{[t]}]=\mathcal{D}[X^{-}_{[t]}]$. Eq.(III) is the so-called integral FT, which tells that the entropy production is non-negative in any macroscopic process, i.e. $\left\langle\Sigma_{\tilde{X}_{[t]}}\right\rangle\geq 0$.
@@
$\Sigma_{x_{[t]}}=\ln\frac{\Pr[\tilde{x}_{[t]}=x_{[t]}]}{\Pr[\tilde{x}^{-}_{[t]}=x^{-}_{[t]}]}$
@@
where $v^{\mu}$ is the velocity of the BP. According to eq. (14), the trajectory heat is

### 28 Feb 2024 v2

@@ I Introduction
Historically, there have been some repeated attempts in the construction of LE on Riemannian manifolds in the mathematical [14, 15, 16, 17, 18, 19] and physical [20, 21] literature. However, the first law and other thermodynamic relations were not considered in these works. Xing et al. [22, 23] explored an Ito-type nonlinear LE and established a FT on Riemannian manifold. Most of these works have interpreted the word “covariance” in the sense of second order (or jet bundle) geometry, which is different from the usual coordinate covariance in standard Riemannian geometry. Ref. [20] seems to be an exception. However, the difference between our work and [20] is acute: the LE in [20] is presented only in the configuration space, something like the OLE to be described in Section II.2. However, as will be explained near the end of Section II.2, the covariance of LE in [20] actually holds only in flat manifolds and only with respect to linear coordinate transformations, while our work covers both the configuration space (OLE) and phase space (ULE) descriptions, and our formalism is genuinely covariant under general coordinate transformation in the sense of first order geometry.
@@ II.1 Underdamped case
In Ref. [24], the relativistic covariant underdamped Langevin equation (ULE) on pseudo-Riemannian spacetime is established. The same procedure can be used for constructing the LE for a point particle of mass $m>0$ moving on $d$-dimensional Riemannian space $M$ with positive definite metric $g_{\mu\nu}(x)$, so we directly present the result,
@@
where $R^{\mu}{}_{a}$ represent the stochastic amplitudes which may depend on $\tilde{x}^{\mu}$ and $\tilde{p}^{\mu}$, $\Gamma^{\mu}{}_{\alpha\beta}$ is the Christoffel connection associated with $g_{\mu\nu}(x)$, $K^{\mu}{}_{\nu}$ is the tensorial damping coefficient (referred to as damping tensor henceforth), and ${\rm d}\tilde{w}^{a}_{t}$ are Gaussian noises with probability distribution functions (PDFs)
@@
The symbol $\circ_{S}$ represents the Stratonovich coupling which ensures the chain rule in stochastic calculus. Greek indices $\mu,\nu,\cdots$ label spatial directions and Latin indices $a,b,\cdots$ label independent noises, which are all running from 1 to $d$. Tilded variables like $\tilde{x}^{\mu},\tilde{p}^{\mu}$ represent random variables, and the un-tilded symbols like $x^{\mu},p^{\mu}$ their realizations. In particular, $(x^{\mu},p^{\mu})$ denotes the coordinate of the Brownian particle on $TM$ in a concrete realization.
@@
(i) The LE described above is a system of stochastic differential equations on the tangent bundle $TM$ regarded as the space of micro states of the Brownian particle. Normally, the space of micro states for a particle is taken to be the cotangent bundle $T^{\ast}M$. However, due to the non-degeneracy of the metric $g_{\mu\nu}$, the tangent and cotangent spaces are dual to each other, and the tangent and cotangent bundles can be used interchangeably (see [25, 26, 27] for explicit use of both approaches). Recall that $TM$ is naturally equipped with the Sasaki metric [25, 26, 28]
@@
As explained in [25, 26], the Sasaki metric is closely connected to the symplectic structure on $TM$.
Notice that the metric $g_{\mu\nu}(x)$ on $M$ plays an indispensable role while obtaining the above Sasaki metric and also while describing the last term of eq. (2).
(ii) The choice for the stochastic amplitude $R^{\mu}{}_{a}$ is non-unique. Different choices correspond to different Langevin systems, and the result of this work should hold for all choices such that $R^{\mu}{}_{a}$ is an invertible matrix function which is differentiable in $(x^{\mu},p^{\mu})$ and transforms as a vector for each fixed $a$.
(iii) Our approach to LE is the traditional one, as opposed to the more abstract nonlinear approach used in [22, 23]. Eq. (2) can be viewed as the geodesic equation supplemented by additional force terms, including a stochastic force
@@
a damping force $f^{\mu}_{\text{dp}}:=-K^{\mu}{}_{\nu}\tilde{p}_{t}^{\nu}/m$ and an external force $f^{\mu}_{\text{ex}}$. The second term in the stochastic force is known as additional stochastic force [29, 30], which is required in order for the Brownian particle to be able to reach thermal equilibrium with the heat reservoir. In one-dimensional case, the stochastic force can also be expressed in the form of post-point rule $\xi=R\circ_{p}{\rm d}\tilde{w}_{t}/{\rm d}t$, hence some authors argued that the post-point rule is better suited for LE with multiplicative noises. However, in higher-dimensional cases, the post-point rule leads to a different result,
@@
(iv) The Stratonovich coupling maintains the chain rule, which ensures the covariance in the usual sense in Riemannian geometry. This makes an important difference from the previous works [16, 17, 18, 19, 22, 23].
(v) Although eqs. (1)-(2) look the same as their relativistic counterparts [24], there are some essential differences. First, the time $t$ used here is absolute, meaning that eqs. (1)-(2) are non-relativistic; Second, the heat reservoir hiding behind the stochastic and damping force terms is also non-relativistic, there is no need to worry about the relativistic effects such as the Tolman-Ehrenfest red shift; Last, the momentum space is flat, as opposed to the relativistic case.
The external force term depends on the position $x$ of the Brownian particle and an external control parameter $\lambda$, and can be separated into conservative and non-conservative parts,
@@
of the Brownian particle. In the extremal case with $\hat{f}^{\mu}=-f^{\mu}_{\text{con}}$, The energy will be consisted purely of the kinematic energy.
@@
where $\tilde{E}_{t}$ is the energy considered as a random variable and $E$ its realization. The part of the increase of energy caused by the heat reservoir is purely a thermal effect and thus comprehended as trajectory heat ${\rm d}\tilde{Q}_{t}$, the rest part is purely mechanical and should be comprehended as trajectory work ${\rm d}\tilde{W}_{t}$,
@@
The ambiguity in the decomposition of the external force also leads to an ambiguity in the trajectory work. However, in any case, the trajectory heat ${\rm d}\tilde{Q}_{t}={\rm d}\tilde{E}_{t}-{\rm d}\tilde{W}_{t}$ is always unambiguously defined. Please be reminded that, unlike the usual heat and work in standard thermodynamics which are inexact differentials defined on the space of macro states, the trajectory heat and work are only defined on a stochastic trajectory of the Brownian particle.
The Fokker-Planck equation (FPE) associated to eqs. (1)-(2) can also be established on $TM$. The PDF for the Brownian particle under the measure ${\rm d}^{2d}X$ is denoted as $\Phi_{t}(x,p):=\Pr[\tilde{x}_{t}=x,\tilde{p}_{t}=p]$, and is clearly coordinate-independent. By use of the diffusion operator method [31], one can get
@@
is the Liouville vector field on $TM$ [25, 26].
Let the non-conservative force be temporarily turned off and the external control parameter be fixed. Then, after sufficiently long period of time, the Brownian particle will reach a thermal equilibrium with the heat reservoir, yielding the equilibrium PDF
$\Phi_{t}(x,p)=\frac{1}{Z}\exp\left[-\frac{1}{T}\left(\frac{g_{\mu\nu}p^{\mu}p^{\nu}}{2m}+U(x,\lambda)\right)\right].$ (7)
@@
$D^{\mu\nu}=2TK^{\mu\nu},$ (8)
which implies that the damping tensor $K^{\mu\nu}$ are not independent of the stochastic amplitudes $R^{\mu}{}_{a}$ and that $K^{\mu\nu}$ is invertible as a matrix. As long as only the FPE is concerned, there is an additional freedom in the sign choice of $R^{\mu}{}_{a}$, because $D^{\mu\nu}$ appear as a quadratic form in $R^{\mu}{}_{a}$.
To facilitate the discussion about FT, we introduce the time-reversal transform (TRT) for the process ranging from $t_{I}$ to $t_{F}$:
$I:\left\{\begin{array}[]{l}x^{\mu}(t)\mapsto x^{\mu}(t_{F}+t_{I}-t)\cr p^{\mu}(t)\mapsto-p^{\mu}(t_{F}+t_{I}-t)\end{array}\right.,$
which is often briefly described as $I:(x,p)\mapsto(x,-p)$ for short. Notice that the infinitesimal time increment ${\rm d}t$ is not affected by such transformation and remains to be positive.
It is obvious that TRT preserves the metric, i.e. $I^{*}\hat{g}=\hat{g}$. The damping force $f^{\mu}_{\text{dp}}|_{X}=-K^{\mu}{}_{\nu}p^{\nu}/m$ reverses sign under TRT, thus the damping tensor must be invariant under TRT,
@@
Eq. (8) implies that the diffusion tensor is also invariant under TRT. Since there is some freedom in choosing the stochastic amplitudes, and hence also in determining their behaviors under TRT. Here we assume the simplest transformation rule,
@@
The last equation is the starting point for introducing the discretized version of the ULE in Appendix B.2.
@@ II.2 Overdamped case
The stochastic mechanics characterized by LE is a branch of physics with multiple time scales, of which the smallest one is the time scale ${\rm d}t$ which allows for a sufficient number of collisions between the Brownian particle and the heat reservoir particles which cause little changes in the state of the Brownian particle [32]. If the temporal resolution $\Delta t$ greatly exceeds ${\rm d}t$ but is still smaller than the relaxation time, the ULE (1)-(2) emerges. If $\Delta t$ is further increased so that it is much larger than the relaxation time, the overdamped limit emerges.
The relaxation process of the Brownian particle can be viewed either as the process by which the damping force attains a state of mechanical equilibrium with other forces (mechanical relaxation process), or as the process by which the Brownian particle achieves local thermal equilibrium with the heat reservoir (thermodynamic relaxation process). The characteristic timescales associated with both processes are $m/\kappa$, where $\kappa$ is the eigenvalue of $K^{\mu\nu}$. Consequently, the OLE should arise when $\Delta t$ greatly exceeds $m/\kappa$.
@@
This result is achieved in flat space under the condition that the damping tensor is position- and momentum-independent [33, 34], and is often viewed as a stochastic differential equation in configuration space. However, if the stochastic amplitudes are momentum-dependent, so is the additional stochastic force. Thus the mechanical equilibrium condition cannot be understood as a stochastic differential equation in configuration space. To avoid the above difficulty, let us consider the simpler situation in which the stochastic amplitudes are momentum-independent. Using the thermodynamic relaxation approach [35], it will be shown that, even in this simpler situation, a nontrivial additional stochastic force term still arises in the corresponding OLE.
@@
$\rho_{t}(x)=\frac{1}{Z_{x}}{\rm e}^{-U(x,\lambda)/T}$
is a solution of eq. (12) provided the non-conservative force is turned off and the external parameter is fixed, wherein $Z_{x}$ represents the configuration space normalization factor, which should not be confused with the normalization factor $Z$ appeared in eq. (7).
@@
$\hat{R}^{\mu}{}_{a}:=\hat{K}^{\mu}{}_{\nu}R^{\nu}{}_{a},\,\hat{F}^{\mu}=\hat{K}^{\mu}{}_{\nu}f^{\nu}_{\text{ex}}+\frac{1}{2}\hat{R}^{\mu}{}_{a}\nabla_{\nu}\hat{R}^{\nu}{}_{a},$ (14)
@@
This equation is similar in form to the LE presented in [21]. However, unlike eq. (14), the additional stochastic force presented in [21] contains only an ordinary coordinate derivative rather than covariant derivative. Consequently, the claimed general covariance of the LE of [20] is questionable: it actually holds only for flat manifolds and only with respects to linear coordinate transformations.
Since the inertial effect can be ignored in the overdamped case, the energy of the Brownian particle contains only the potential energy, i.e. $E(x,\lambda)=U(x,\lambda)$. Using eq. (10) and the chain rule, we have
@@ III Fluctuation theorem
Now we come to the stage for describing FT on Riemannian manifold based on the description of stochastic trajectories. Since the trajectory probability in continuous time is hard to deal with, we adopt the following strategy: first we take a discrete equidistant set of time nodes $t_{I}=t_{0}<t_{1}<...<t_{n}=t_{F}$ to rewrite the LE and factorize the corresponding trajectory probabilities, and then take the continuum limit at the end of the calculation.
@@
Before proceeding, it is necessary to clarify the concepts of ensemble and trajectory entropy productions. At the time $t$, the ensemble entropy of the Brownian particle reads
@@
[…]
@@
is not the entropy of the overdamped Brownian particle, however, the ensemble entropy production of the Brownian particle can be represented as the difference of $-\overline{\ln\rho_{t}}$, because the ensemble entropy for the overdamped Brownian particle can be evaluated to be
@@
where the last two terms arise from the momentum space integration of the term involving the distribution $P^{s}(x,p)$ given in eq. (11). Subtracting the initial value from the final value leaves only the difference of $-\overline{\ln\rho_{t}}$,
@@
Since the Brownian motion is a Markov process, the trajectory probability can be decomposed into product of transition probabilities,
$\Pr[\tilde{X}_{[t]}=X_{[t]}]$
$\quad={\color[rgb]{0,0,0}\Big{(}}\prod_{i=0}^{n-1}\Pr[\tilde{X}_{i+1}=X_{i+1}|\tilde{X}_{i}=X_{i}]{\color[rgb]{0,0,0}\Big{)}}\Pr[\tilde{X}_{0}=X_{0}].$
@@
where, since the heat reservoir maintains in equilibrium, the Clausius equality holds, the change of the entropy of the heat reservoir reads
@@
${\rm e}^{-\big{\langle}\Sigma_{\tilde{X}_{[t]}}\big{\rangle}}$ $\leq\big{\langle}{\rm e}^{-\Sigma_{\tilde{X}_{[t]}}}\big{\rangle}=\int\mathcal{D}[X_{[t]}]\Pr[\tilde{X}^{-}_{[t]}=X^{-}_{[t]}]$
@@
where $\mathcal{D}[X_{[t]}]={\rm d}X_{0}\wedge{\rm d}X_{1}\wedge...\wedge{\rm d}X_{n}$ is the measure on the trajectory space. It is easy to prove that the map $X_{[t]}\mapsto X^{-}_{[t]}$ preserves the measure, i.e. $\mathcal{D}[X_{[t]}]=\mathcal{D}[X^{-}_{[t]}]$. Eq.(III) is the so-called integral FT, which tells that the entropy production is non-negative in any macroscopic process, i.e. $\big{\langle}\Sigma_{\tilde{X}_{[t]}}\big{\rangle}\geq 0$.
@@
$\Sigma_{x_{[t]}}{\color[rgb]{0,0,0}:=}\ln\frac{\Pr[\tilde{x}_{[t]}=x_{[t]}]}{\Pr[\tilde{x}^{-}_{[t]}=x^{-}_{[t]}]}$
@@
where $v^{\mu}$ is the velocity of the Brownian particle. According to eq. (16), the trajectory heat is